Circular No. 6845

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

PSR 1744–1134 W. Becker and J. Trümper, Max-Planck-Institut für Extraterrestrische

Physik; and G. Hasinger, Astrophysikalisches Institut, Potsdam, communicate: “We report the detection of the isolated galactic millisecond pulsar PSR 1744–1134 in soft x-rays with ROSAT. The pulsar was observed with the HRI at the focus of the XRT during 1997 Sept. 3–24 for an effective exposure time of 61 238 s. The pulsar’s net countrate (including background, vignetting, and dead-time corrections) is (2.9 ± 0.9) × 10^{−4} count/s. Assuming a Crab-like spectrum and a column absorption of N_{H} = 10^{20} cm^{−2}, deduced from the radio dispersion measure (cf. Bailes et al. 1997, Ap.J. 481, 386), we find an energy flux of 2 × 10^{−14} erg s^{−1} cm^{−2} over 0.1–2.4 keV. The low number of recorded source counts precluded any modulation testing at the 4.07-ms radio period.”

SUPERNOVA 1998S IN NGC 3877 P. Garnavich, R. Kirshner, and P. Challis, Harvard-Smithsonian Cen-

ter for Astrophysics, on behalf of the Supernova Intensive Study team, report that the Space Telescope Imaging Spectrograph aboard the Hubble Space Telescope obtained spectra (range 114–570 nm, via three lowresolution gratings) of SN 1998S on Mar. 16.1 UT. Longward of 300 nm, the spectrum shows primarily Balmer emission with broad, shallow P-Cyg absorption troughs. The maximum velocity measured in the Hβ absorption is −10 000 km/s. The continuum peaks at 265 nm, and the flux estimated from the continuum around 440 nm corresponds to magnitude B = 12.3. Shortward of 300 nm, the character of the spectrum changes radically. Deep, blended absorption lines dominate the spectrum. The strongest lines show a narrow P-Cyg emission (500 km/s FWHM), narrow absorption shifted by −1100 km/s from the emission, and a broad, shallow absorption extending to −10 000 km/s. Major lines include Mg II (280 nm), C III (230 nm), C IV (154 nm), Si IV (140 nm), C II (134 nm), O I (130 nm) and N V (124 nm), Lyα (122 nm), and C III (118 nm). Spectra taken with the Fred L. Whipple Observatory 1.5-m telescope on Mar. 19.5 by D. Koranyi and M. Calkins show Hα with a profile similar to the ultraviolet lines, but with the narrow absorption at a blue shift of only −415 km/s. The broad absorption extends to −9000 km/s, with a minimum at −4700 km/s. Hα has a redshifted emission tail extending to ∼ +1600 km/s.

1998 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green